Tshepo want to buy a big screen TV.She has five interests rates to choose from if she borrows the money from the bank.the cheapest option

*29%per year,compounded daily

*30%per year, compounded semi annually

*28.5%pet year, compounded weekly

*29.5%per year, compounded every 2 months.


















1
Expert's answer
2022-02-23T13:55:39-0500

"FV=PV (1+\frac{r}{m})^{nm}"



Where FV is the future amount to be repaid

PV is the amount borrowed now

r is the rate of interest

m is the number of times compounded

n is the number of periods in years.


Assume PV=1000


A) "FV=1000 (1+\frac{0.29}{365})^{365×5}"

"FV=4260.66"


B) "FV=1000(1+\frac{0.3}{2})^{2×5}"

"FV=4045.56"


C)"FV=1000 (1+\frac{0.285}{52})^{52×5}"

"FV=4141.71"


D) "FV=1000 (1+\frac{0.295}{6})^{6×5}"

"FV=4220.21"


E)"FV=1000 (1+\frac{0.29}{12})^{12×5}"

"FV=4190.23"


Of all these options available, only option B) is the cheapest or lowest in terms of repayment. So she'd opt for 30% per annum compounded semiannually